# GoalFlow: Goal-Driven Flow Matching for Multimodal Trajectories Generation in End-to-End Autonomous Driving

### v4

@@ 4 Experiments
Method Ego Stat. Image LiDAR Video $\mathbf{S_{NC}}$uparrow $\mathbf{S_{DAC}}$uparrow $\mathbf{S_{TTC}}$uparrow $\mathbf{S_{CF}}$uparrow $\mathbf{S_{EP}}$uparrow $\mathbf{S_{PDM}}$uparrow
@@
[…]
Model Description $\mathbf{S_{NC}}$uparrow $\mathbf{S_{DAC}}$uparrow $\mathbf{S_{TTC}}$uparrow $\mathbf{S_{CF}}$ $\mathbf{S_{EP}}$uparrow $\mathbf{S_{PDM}}$uparrow
@@
$T$ $\mathbf{Inf.Time}$ $\mathbf{S_{NC}}$uparrow $\mathbf{S_{DAC}}$uparrow $\mathbf{S_{TTC}}$uparrow $\mathbf{S_{CF}}$uparrow $\mathbf{S_{EP}}$uparrow $\mathbf{S_{PDM}}$uparrow
@@
$\sigma$ $\mathbf{S_{NC}}$uparrow $\mathbf{S_{DAC}}$uparrow $\mathbf{S_{TTC}}$uparrow $\mathbf{S_{CF}}$uparrow $\mathbf{S_{EP}}$uparrow $\mathbf{S_{PDM}}$uparrow
@@ 4.5 Results and Analysis
Dim Backbone $\mathbf{S_{NC}}$uparrow $\mathbf{S_{DAC}}$uparrow $\mathbf{S_{TTC}}$uparrow $\mathbf{S_{EP}}$uparrow $\mathbf{S_{PDM}}$uparrow
256 V2-99 97.1 96.2 91.8 81.8 86.5
512 V2-99 97.3 97.6 92.5 83.0 88.1
1024 V2-99 98.6 97.5 94.7 85.0 89.4
256 resnet34 98.3 93.8 94.3 79.8 85.7
256 V2-99 97.1 96.2 91.8 81.8 86.5
Table 5: Impact of Scaling Model. We examine the impact of scaling the Transformer’s hidden dimension and changing the image backbone. Both increasing the hidden dimension and using a larger backbone improve performance. To isolate the effect of trajectory post-processing, we compare with $\mathcal{M}_{2}$.
Dim Backbone $\mathbf{S_{NC}}$uparrow $\mathbf{S_{DAC}}$uparrow $\mathbf{S_{TTC}}$uparrow $\mathbf{S_{EP}}$uparrow $\mathbf{S_{PDM}}$uparrow

### v6

@@ 4 Experiments
Method Ego Stat. Image LiDAR Video $\mathbf{S_{NC}}$ $\uparrow$ $\mathbf{S_{DAC}}$ $\uparrow$ $\mathbf{S_{TTC}}$ $\uparrow$ $\mathbf{S_{CF}}$ $\uparrow$ $\mathbf{S_{EP}}$ $\uparrow$ $\mathbf{S_{PDM}}$ $\uparrow$
@@
Table 1: Comparisons with SOTA methods in PDM score metrics on Navsim [7] Test. Our method outperforms other approaches across all evaluation metrics. $\dagger$} uses the endpoint of the ground-truth trajectory as the goal point. $\ddagger$} uses the ground-truth trajectories to evaluate.
Model Description $\mathbf{S_{NC}}$ $\uparrow$ $\mathbf{S_{DAC}}$ $\uparrow$ $\mathbf{S_{TTC}}$ $\uparrow$ $\mathbf{S_{CF}}$ $\mathbf{S_{EP}}$ $\uparrow$ $\mathbf{S_{PDM}}$ $\uparrow$
@@
$T$ $\mathbf{Inf.Time}$ $\mathbf{S_{NC}}$ $\uparrow$ $\mathbf{S_{DAC}}$ $\uparrow$ $\mathbf{S_{TTC}}$ $\uparrow$ $\mathbf{S_{CF}}$ $\uparrow$ $\mathbf{S_{EP}}$ $\uparrow$ $\mathbf{S_{PDM}}$ $\uparrow$
@@
$\sigma$ $\mathbf{S_{NC}}$ $\uparrow$ $\mathbf{S_{DAC}}$ $\uparrow$ $\mathbf{S_{TTC}}$ $\uparrow$ $\mathbf{S_{CF}}$ $\uparrow$ $\mathbf{S_{EP}}$ $\uparrow$ $\mathbf{S_{PDM}}$ $\uparrow$
@@ 4.5 Results and Analysis
Dim Backbone $\mathbf{S_{NC}}$ $\uparrow$ $\mathbf{S_{DAC}}$ $\uparrow$ $\mathbf{S_{TTC}}$ $\uparrow$ $\mathbf{S_{EP}}$ $\uparrow$ $\mathbf{S_{PDM}}$ $\uparrow$